OpenStudy (anonymous):

A bowl contains 8 blue marbles and 5 green marbles. Hannah randomly draws 4 marbles from the bowl. She does not replace the marbles after each draw. What is the probability that she draws 2 blue marbles and 2 green marbles?

OpenStudy (misssunshinexxoxo):

8C2*5C2 / 13C4 = 28*10 / 715 = 56/143

OpenStudy (anonymous):

looks good to me did you understand any of it?